OpenStudy (anonymous):

is the set of functions linearly independent or linearly dependent. f1(x)= exp(x+2) , f2(x)=exp(x-3)

OpenStudy (anonymous):

\[f1(x)=\exp(x+2) , f2(x)=\exp(x-3) on (-\infty, \infty)\]